OpenStudy (anonymous):

A particle moves according to a law of motion s = f(t) = t^3 - 9t^2 + 15t, t >= 0, where t is measured in seconds and s in feet. a)Find the total distance traveled during the first 7 seconds

OpenStudy (anonymous):

that is wrong...

OpenStudy (anonymous):

F(t)= position F(t)= Velocity=\[(3 t^2 - 18 t + 15)^2\] Speed=\[\sqrt{\left(3 t^2-18 t+15\right)^2}\] Distance=\[\int_0^7 \sqrt{\left(3 t^2-18 t+15\right)^2} \, dt\]=71

OpenStudy (anonymous):

WOW it was right....Thank you

OpenStudy (anonymous):

I knew it was not gonna be easy as plugging a number in

OpenStudy (anonymous):

but why will you square something under a square root

OpenStudy (anonymous):

i never seen this before

OpenStudy (anonymous):

to take the absolute value